### v1

@@ 1 Introduction
Early researches [67, 18, 48, 56, 24, 37] converted point cloud data into multi-view[59, 6, 5, 4] or voxel representations, utilizing traditional convolutional neural networks to learn shape information. However, this conversion process often
@@
led to the loss of inherent geometric information and incurred high computational costs. PointNet [29] directly encoded each point in the point cloud independently and aggregated global features through max pooling, but this approach overlooked local structural information. To address this issue, subsequent works [47, 46, 17, 35] proposed a series of methods based on local feature aggregation. These methods divide the point cloud into different local subsets through farthest point sampling (FPS), then learn local shape representations by constructing local aggregation operators, and finally learn from local to global shape perception by constructing a hierarchical structure. However, such methods overlook the long-range dependency relationships among points.
@@ 3.3 Local Structure Fitting Convolution
3.3.1 Taylor Series.
Inspired by Taylor series (see Fig. 2(a)), building upon GPM, we adopt a local fitting approach to analyze the local structure and details of the point cloud more finely. The Taylor series is given by:
@@
$f(x)\approx f(a)+\sum_{n=1}^{\infty}a_{n}(x-a)^{n},\quad|x-a|<\epsilon,$ (10)
@@
3.3.2 Local Structure Fitting Convolution.
@@
where we refer to $f_{i}^{L}$ as Low-Order Convolution (LOConv) and $f_{i}^{H}$ as High-Order Convolution (HOConv, see Fig. 2(b)). $\phi$ represents MLP. $\mathcal{T}(f_{i},f_{j})$ is a novel affine basis function. Here, ||$\cdot$|| ⋅ | denotes element-wise absolute value, $s\in\{0,1\}$, and $p$ is a learnable parameter. When $s=1,p=1$, $\mathcal{T}(f_{i},f_{j})$ degenerates into an Affine Basis Function (ABF [38]) (see Eq. 13); when $s=0,p=2$, $\mathcal{T}(f_{i},f_{j})$ degenerates into a radial basis function (RBF [25]) (see Eq. 14). Therefore, the proposed $\mathcal{T}$ exhibits powerful representation capability.
@@
3.3.3 Explicit Structure Introduction.
@@ 5 Conclusion
In this paper, we propose a novel Global Perception and Local Structure Fitting-based Transformer (GPSFormer) to address the challenge of effectively capturing shape information from irregular point clouds. The key contributions of GPSFormer include introducing a Global Perception Module (GPM) and a Local Structure Fitting Convolution (LSFConv). The GPM enhances the model’s ability to capture global contextual information in point clouds by introducing the Adaptive Deformable Graph Convolution (ADGConv). Meanwhile, the LSFConv finely learns the local geometric structures of point clouds, acquiring both low-order fundamental information and high-order refinement details. Through extensive experiments in three point cloud understanding tasks, the proposed GPSFormer demonstrates efficient processing and analysis of point clouds without relying on external data. In the future, we plan to further explore the potential of GPSFormer in pre-training, lightweight approaches, and few-shot learning settings.

### v2

@@ 1 Introduction
Early researches [67, 18, 48, 56, 24, 37] converted point cloud data into multi-view[59, 6, 5, 4] or voxel representations, utilizing traditional convolutional neural networks to learn shape information. However, this conversion process often led to the loss
@@
of inherent geometric information and incurred high computational costs. PointNet [29] directly encoded each point in the point cloud independently and aggregated global features through max pooling, but this approach overlooked local structural information. To address this issue, subsequent works [47, 46, 17, 35] proposed a series of methods based on local feature aggregation. These methods divide the point cloud into different local subsets through farthest point sampling (FPS), then learn local shape representations by constructing local aggregation operators, and finally learn from local to global shape perception by constructing a hierarchical structure. However, such methods overlook the long-range dependency relationships among points.
@@ 3.3 Local Structure Fitting Convolution
3.3.1 Taylor Series
Inspired by Taylor series, building upon GPM, we adopt a local fitting approach to analyze the local structure and details of the point cloud more finely. The Taylor series is given by:
@@ 3.3.1 Taylor Series
$f(x)=f(a)+\sum_{n=1}^{\infty}a_{n}(x-a)^{n},\quad|x-a|<\epsilon,$ (10)
@@
3.3.2 Local Structure Fitting Convolution
@@
where we refer to $f_{i}^{L}$ as Low-Order Convolution (LOConv) and $f_{i}^{H}$ as High-Order Convolution (HOConv). $\phi$ represents MLP. $\mathcal{T}(f_{i},f_{j})$ is a novel affine basis function. Here, ||$\cdot$|| ⋅ | denotes element-wise absolute value, $s\in\{0,1\}$, and $p$ is a learnable parameter. When $s=1,p=1$, $\mathcal{T}(f_{i},f_{j})$ degenerates into an Affine Basis Function (ABF [38]) (see Eq. 13); when $s=0,p=2$, $\mathcal{T}(f_{i},f_{j})$ degenerates into a radial basis function (RBF [25]) (see Eq. 14). Therefore, the proposed $\mathcal{T}$ exhibits powerful representation capability.
@@
3.3.3 Explicit Structure Introduction:
@@ 5 Conclusion
In this paper, we propose a novel Global Perception and Local Structure Fitting-based Transformer (GPSFormer) to address the challenge of effectively capturing shape information from irregular point clouds. The key contributions of GPSFormer include introducing a Global Perception Module (GPM) and a Local Structure Fitting Convolution (LSFConv). The GPM enhances the model’s ability to capture global contextual information in point clouds by introducing the Adaptive Deformable Graph Convolution (ADGConv). Meanwhile, the LSFConv finely learns the local geometric structures of point clouds, acquiring both low-order fundamental information and high-order refinement details. Through extensive experiments, the proposed GPSFormer demonstrates efficient processing and analysis of point clouds without relying on external data. In the future, we plan to further explore the potential of GPSFormer in pre-training, lightweight approaches, and few-shot learning settings.
Acknowledgment
This work was supported in part by NTU-DESAY SV Research Program under Grant 2018-0980; and in part by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1, under Grant RG78/21. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).